OpenStudy (yamyam70):

rewrite as single logarithm 2 log u - 3 log v - 2 log z

OpenStudy (yamyam70):

Is it...

OpenStudy (yamyam70):

\[\log \frac{ u^2 }{v^3z^2 } or \log \frac{ \frac{ u^2 }{ v^3 } }{ z^2 } ? \]

OpenStudy (mathstudent55):

Those two are the same.

OpenStudy (yamyam70):

Thanks man :)

OpenStudy (mathstudent55):

wlcm